OpenStudy (anonymous):

Write the rate as a unit rate. Mickael Saltwood can paint 138 doors in 10 hours, while Scott Kepler can paint 59 doors in 5 hours. Who can paint doors faster, Mickael or Scott?

OpenStudy (anonymous):

59x2=118 so Mickael is faster

OpenStudy (anonymous):

oh unit rate sorry 10*60 to get the minutes 600/138doors = 4.35minutes to paint one door for Mickeal 5*60 = 300 300/59 = 5.08minutes for Scott to paint a door. aka Mickael is faster. answere any good ?